The solubility of Ag2CrO4 at 25 °C in water is 0.0027 g/100 mL. Estimate Ksp for this substance?
Answer #1

Solubility = 0:0027g|ioome = 2-7 X 10 ²g/100ml = 2.7 x 103 1011 = 2.7 X 10² g/L . Molar solubility = 27x10² g/L 331.7296 g